How would you do this problem...2x^2+5x+3=0?
Found 2 solutions by solver91311, edjones:
Answer by solver91311(24713)



I would either factor it, complete the square, or use the quadratic formula.


John


Answer by edjones(8007)

2x^2+5x+3=0
2x^2+2x+3x+3=0
2x(x+1)+3(x+1)=0
(2x+3)(x+1)=0
.
2x+3=0
2x=-3
x=-3/2
.
x+1=0
x=-1
.
Ed